OpenStudy (anonymous):

Can someone help with this equation T/3 - 2 = 1/3

OpenStudy (anonymous):

T=7

OpenStudy (anonymous):

T/3-7/3 = 0

OpenStudy (anonymous):

(T-6)/3 = 1/3

OpenStudy (anonymous):

T=7

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{T }{3} -2=\frac{1}{3} }\) \(\large\color{black}{ \displaystyle \frac{T }{3} -\frac{6}{3}=\frac{1}{3} }\) \(\large\color{black}{ \displaystyle T-6=1 }\) \(\large\color{black}{ \displaystyle T=? }\)